### v1

# A multiple imputation approach to distinguish curative from life-prolonging effects in the presence of missing covariates

Abstract
Medical advancements have increased cancer survival rates and the possibility of finding a cure. Hence, it is crucial to evaluate the impact of treatments in terms of both curing the disease and prolonging survival. We may use a Cox proportional hazards (PH) cure model to achieve this. However, a significant challenge in applying such a model is the potential presence of partially observed covariates in the data. We aim to refine the methods for imputing partially observed covariates based on multiple imputation and fully conditional specification (FCS) approaches. To be more specific, we consider a more general case, where different covariate vectors are used to model the cure probability and the survival of patients who are not cured. We also propose an approximation of the exact conditional distribution using a regression approach, which helps draw imputed values at a lower computational cost. To assess its effectiveness, we compare the proposed approach with a complete case analysis and an analysis without any missing covariates. We discuss the application of these techniques to a real-world dataset from the BO06 clinical trial on osteosarcoma.
Keywords mixture cure models  $\cdot$ multiple imputation  $\cdot$ chained equations  $\cdot$ osteosarcoma
1 Introduction
Recent advancements in medicine have increased the chances of survival for some types of cancer such as breast cancer, melanoma, osteosarcoma, and childhood acute lymphocytic leukemia, and have even made cure possible. As a result, it is important not only to estimate the impact several risk factors may have on survival but also to analyze their effect on the probability of being cured. This is especially relevant in pediatric oncology. Osteosarcoma, a type of bone tumour mainly affecting children and young adults, may help illustrate this need. For this type of cancer, over 95% of patients who do not experience a relapse within 5 years after treatment are considered cured and are unlikely to relapse later (Ferrari et al., 2006). Hence, it is crucial to consider cured patients in statistical analyses and to assess the efficacy of a treatment in terms of cure rather than only focusing on its impact on prolonging survival.
In survival analysis, individuals who will never experience the event of interest, such as disease recurrence or disease-related death, are referred to as "cured". Traditional survival analysis methods can be less informative when a portion of the population is considered cured, as they may not account for differences in cure status within the population. Additionally, the cure status is not always observable; while we know that patients who experience the event of interest are uncured, we cannot determine if censored patients will eventually experience the event. This led to the development of cure models, which provide a more suitable modelling approach, particularly in oncological studies (Legrand and Bertrand, 2019).
One type of cure model is the mixture cure model, which defines overall survival as a combination of the cure probability (incidence) and the survival for uncured patients (latency). Mixture cure models allow for the use of different covariates when modelling the incidence and the latency components. This means we can differentiate between the effect of a covariate on the probability of being cured and its effect on the survival of uncured patients. In this context, we consider a Cox proportional hazards (PH) cure model, where the incidence is modelled by a logistic regression model and a Cox PH model is employed for the latency. Estimation is generally carried out using a maximum likelihood approach in an incomplete data setting. For this purpose, the expectation-maximization (EM) algorithm is often used due to the latent cure status (Cai et al., 2012). For more detailed information on cure models, see the review paper by Amico and Van Keilegom (Amico and Van Keilegom, 2018) or the book of Peng and Yu (Peng and Yu, 2021).
A recent study by Musta et al. (Musta et al., 2022) examined the prognostic significance of histologic response and intensified chemotherapy on the cure rate and progression-free survival among osteosarcoma patients who have not been cured. However, the study excluded observations with missing histological responses from the analysis. In this article, we propose a method based on a multiple imputation approach, to analyze a "completed" dataset.
When working with an incomplete dataset, a simple but naive approach to handling missing data is to conduct a complete-case analysis, which involves removing observations with missing data. However, this method can lead to inefficient and inconsistent results when the missing mechanism is not completely at random (Little and Rubin, 2002). Alternatively, missing values can be filled in by using simple imputation techniques, either based on summaries (e.g. mean) or statistical models (e.g. regression) based on available data. While these methods allow for the use of the whole sample, completed by imputation, they may underestimate the final model’s standard errors (van Buuren, 2018). To address these issues, Rubin (Rubin, 1987) introduced multiple imputation techniques. This approach produces unbiased estimates and incorporates the uncertainty of imputation into standard error estimates. Multiple imputation involves creating multiple imputed datasets, estimating parameters in each of them separately, and then combining the results. When dealing with survival data, it has been found beneficial to include the survival outcome in the imputation model for the covariates (White and Royston, 2009).
However, when dealing with cured patients, the latent cure status presents an additional challenge for the imputation procedure. Beesley et al. (Beesley et al., 2016) have developed a method for using multiple imputation in the Cox proportional hazards (PH), which is based on specifying an exact or approximated distribution of the missing covariates, conditional on the available ones and the survival response (time, status). However, this method assumes that the same design vector is used in the models for both the incidence and latency of the cure, which can be a limitation as including all available covariates in both components of the mixture model may lead to the estimation of a large number of parameters. Additionally, the factors affecting the probability of being cured may differ from those influencing the hazard, making the settings of incidence and latency distinct from each other. Lastly, as demonstrated by Hanin and Huang (Hanina and Huang, 2014), sharing covariates between various components of a Cox PH cure model may prevent the model identifiability. Hence, a simpler model would be preferred (over the one with all the covariates in both components) whenever supported by the data.
In this paper, we propose to release this constraint and to generalize the existing methodology for multiple imputation in cure models by accommodating possibly different sets of covariates when modelling the cure probability and the survival of the uncured patients. This allows us to fully leverage the benefits of a mixture model. Our motivation for this improvement also stems from its application to the analysis of osteosarcoma data (Musta et al., 2022). The results of our simulation study also emphasise the need to differentiate between the covariates modelling the cure probability and the survival of the uncured patients. Particularly, we observe that, when a covariate with missing values is wrongly included in both models, multiple imputation based on the approximate method performs better than the exact conditional distribution approach.
In Section 2, we introduce the fundamental concepts underlying cure models and the use of multiple imputation in these models. In Section 3, we outline the methodology for performing multiple imputation in cure models when the sets of covariates for incidence and latency models are different. In Section 4, we evaluate the performance of this method using simulated data. In Section 5, we apply the developed methodology to analyze osteosarcoma data from the BO06 clinical trial and compare the results with the complete-case approach for sensitivity. Finally, in Section 6, we conclude with a discussion.
2 Basic concepts
2.1 Mixture cure models
We are interested in studying the time $T$ until an event of interest occurs, in a sample consisting of both cured and uncured individuals. The cured individuals will not experience the event of interest during their lifetime. We denote by $G$ the binary indicator for the uncured status, i.e. $G=1$ if uncured and $G=0$ if cured. However, due to censoring at time $C$ and to the limited time window covered by the study at hand, these variables can not be directly observed. Instead, we observe the follow-up time $Y=\min(T,C)$ and the event indicator $\Delta=I(T<C)$. For uncensored observations ($\Delta=1$), $G=0$, while for censored observations $G$ is unknown.
We consider a mixture cure model, which expresses overall survival as a combination of the probability of being cured (with potentially infinite survival) and the survival of the uncured individuals. The probability of being uncured, also known as incidence, given a design vector $X\in\mathbb{R}^{p}$, is denoted by $\pi(X)=P(G=1|X)$, and is typically described by a logistic regression function (Legrand and Bertrand, 2019; Farewell, 1977)
$\pi(X)=\frac{\exp{(\alpha_{0}+\alpha^{T}X)}}{1+\exp{(\alpha_{0}+\alpha^{T}X)}}$ (1)
where $\alpha_{0}\in\mathbb{R}$ is the intercept and $\alpha\in\mathbb{R}^{p}$ denotes the vector of regression parameters.
The survival of the uncured patients, known as latency, given a set of covariates $Z\in\mathbb{R}^{q}$, which may differ from $X$, is represented by $S_{u}(t|Z)=P(T>t|Z,G=1)$, and it can be modelled employing a regression model such as the Cox PH model (Cox, 1972)
$S_{u}(t|Z)=\exp\left(-H_{0}(t)\exp{(\beta^{T}Z)}\right)$ (2)
where $H_{0}$ denotes the cumulative baseline hazard function of the uncured and $\beta\in\mathbb{R}^{q}$ is the vector of regression parameters. In this model, the proportional hazards assumption holds for the sub-population of uncured patients, but not for the entire population. Combining these components, see (Kuk and Chen, 1992), the overall survival is given by
$S(t|X,Z)=(1-\pi(X))+\pi(X)S_{u}(t|Z)$ (3)
This model is referred to as the Cox PH cure model.
For a cure model to be identifiable, a sufficiently long follow-up is required, i.e. the study duration is longer than the support of the distribution for the event times of the uncured individuals. For a more extensive discussion regarding the assumption of sufficient follow-up, we refer the reader to the review by Amico and Van Keilegom (Amico and Van Keilegom, 2018). In practice, to check whether such an assumption is satisfied, one typically combines medical knowledge with a visual inspection of the Kaplan-Meier estimator. Ideally, we would like to observe a long plateau after the last observed event, containing many censored observations. Recently, some statistical tests have also been proposed for the sufficient follow-up assumption (Xie et al., 2023; Yuen and Musta, 2024; Maller et al., 2024).
The complete-data likelihood function for the Cox PH cure model can be written as
$\mathcal{L}_{c}(\alpha_{0},\alpha,\beta,H_{0})=\prod_{i=1}^{n}[\pi(X_{i})f_{u}(Y_{i}|Z_{i})]^{\Delta_{i}G_{i}}[\pi(X_{i})S_{u}(Y_{i}|Z_{i})]^{(1-\Delta_{i})G_{i}}[1-\pi(X_{i})]^{(1-\Delta_{i})(1-G_{i})}$ (4)
and it can be factorized into two components, containing the parameters of the incidence and the latency, respectively.
Since this model depends on the latent cure status $G$, the set of parameters is usually estimated using an Expectation-Maximization (EM) algorithm (Peng and Dear, 2000; Sy and Taylor, 2000), which is an iterative two-step procedure. In the first step, the posterior expectation of the complete data log-likelihood over the latent uncured indicator is obtained, conditional on the parameters estimated from the previous iteration. For individual $i\in(1,...n)$ at iteration $r$, this is given by
$q_{i}^{(r)}$ $=E(G_{i}|Y_{i},\Delta_{i},X_{i},Z_{i},\alpha_{0}^{(r-1)},\alpha^{(r-1)},\beta^{(r-1)},H_{0}^{(r-1)})$
$=\Delta_{i}+(1-\Delta_{i})\frac{\pi^{(r-1)}(X_{i})S_{u}^{(r-1)}(Y_{i}|Z_{i})}{1-\pi^{(r-1)}(X_{i})+\pi^{(r-1)}(X_{i})S_{u}^{(r-1)}(Y_{i}|Z_{i})}$
where $\pi^{(r-1)}$ and $S_{u}^{(r-1)}$ are computed according to the equations (1) and (2) by using the parameter estimates resulting from iteration $r-1$, namely $\alpha_{0}^{(r-1)},\alpha^{(r-1)},\beta^{(r-1)},H_{0}^{(r-1)}$. The second step of the EM algorithm consists of maximizing the expected log-likelihood for complete data given in (4), where $G_{i}$ is replaced by its expectation $q_{i}^{(r)}$ in each iteration. This reduces to a simpler weighted data problem, with an expected log-likelihood which can be maximized by standard algorithms.
If parameter distinctiveness holds, the two components of the expected likelihood can be maximized separately, one with respect to $\alpha_{0}$ and $\alpha$, and the other with respect to $\beta$ and $H_{0}$. Maximization with respect to $H_{0}$ at iteration $r$ leads to the following modified Breslow estimator (Breslow, 1974) of the baseline cumulative hazard function
$\hat{H}_{0}^{(r)}(t)={\sum\limits_{t_{j}\leq t}}\frac{D(t_{j})}{{\sum\limits_{i\in R_{j}}}q_{i}^{(r)}\exp{(\beta^{(r-1)T}Z_{i})}}$ (5)
where $D(t_{j})$ is the number of events at time $t_{j}$ and $R_{j}$ represents the set of observations at risk at time $t_{j}$ (Sy and Taylor, 2000).
2.2 Handling missing values by multiple imputation
We first illustrate the multiple imputation approach for a general model with outcome $Y$ and design vector $X\in\mathbb{R}^{p}$, where $X^{(j)}$ has missing observations. The procedure consists of imputing the incomplete covariate $X^{(j)}$ defining a model for this element, conditional on the outcome $Y$ and all the remaining covariates $X^{(-j)}=X\setminus X^{(j)}$. This is done through the fully conditional specification (FCS) (van Buuren et al., 2006) approach, requiring the specification of the conditional distribution $f(X^{(j)}|Y,X^{(-j)};\Psi^{j}$), where $\Psi^{j}$ represents a set of parameters. The procedure begins by fitting a regression model for $X^{(j)}$ with covariates $Y$ and $X^{(-j)}$ on the portion of observed data for $X^{(j)}$. This provides us with estimates $\hat{\Psi}^{(j)}$ for the regression coefficients $\Psi\in\mathbb{R}^{p}$, as well as the estimated covariance matrix $\hat{\Sigma}_{\Psi}$. We then draw values $\Psi^{*}$ from the posterior distribution of $\Psi$ given the observations, which is frequently assumed to be $N(\hat{\Psi}^{(j)},\hat{\Sigma}_{\Psi})$. Imputed values for the cases where $X^{(j)}$ is missing are drawn using this conditional distribution $f(X^{(j)}|Y,X^{(-j)};\hat{\Psi}^{*})$. This step is repeated $K$ times, resulting in $K$ imputed samples.
When there are multiple variables with missing observations, this process is repeated sequentially for each one of them, using the observed and the imputed values of the remaining variables. This procedure, known as multiple imputation by chained equations, is summarized in Figure 1. It starts by filling all missing values randomly and it then creates $K$ datasets with imputed values for the incomplete variables. The order in which the variables are imputed is determined by the proportion of missing cases, starting with the variable that has the least missing data. (van Buuren, 2018). Each variable uses the observed and the most recently imputed values of the remaining ones, and this method is iterated on each dataset until convergence.
Figure 1: Multiple imputation by chained equations procedure
After obtaining $K$ imputed datasets, the set of parameters $\Psi$ is estimated separately using each of these datasets, resulting in a vector of estimates $(\hat{\Psi}_{1},...,\hat{\Psi}_{K})$ and estimated covariance matrices $(\hat{\Sigma}_{1},...,\hat{\Sigma}_{K})$. Finally, the pooled estimate is obtained by combining the $K$ estimates using Rubin’s rules (Rubin, 2004)
$\hat{\Psi}=\frac{1}{K}\sum\limits_{k=1}^{K}\Psi_{k}$
and the final covariance matrix is given by
$\hat{\Sigma}_{\Psi}=\frac{1}{K}\sum\limits_{k=1}^{K}\Sigma_{k}+\left(1+\frac{1}{K}\right)\frac{1}{K-1}\sum\limits_{k=1}^{K}(\Psi_{k}-\hat{\Psi})(\Psi_{k}-\hat{\Psi})^{T}$
where the first component is the average of the repeated complete-data posterior covariances of $\Psi$ (within component), and the second component is the estimate of the covariance between the $K$ complete-data posterior means of $\Psi$ (between component).
2.3 Multiple imputation for cure models
To perform multiple imputation in the case of a cure model with missing covariates, the first step is to determine the conditional distribution of a covariate with missing values given the outcome and the other covariates. For survival data, the outcome consists of both the follow-up time $Y$ and the event indicator $\Delta$. In the case of the standard Cox model (without cure fraction), it has been shown that a suitable model for imputing the missing covariate values is based on $\Delta$, the cumulative baseline hazard $H_{0}(Y)$ and the other covariates (White and Royston, 2009).
However, for the cure model, an additional challenge arises from the fact that even the cure status $G$ is unobserved for the censored observations. For a Cox PH cure model where the same design vector is common to both the incidence and the latency (i.e. $X=Z$), Beesley et al. (Beesley et al., 2016) derived the exact conditional distribution based on the complete likelihood function in (4). They proposed a chained equations procedure which imputes iteratively the missing values for the cure status $G$ and the incomplete covariates. In practice, drawing observations from the exact conditional distribution, which has a complicated form, is often computationally expensive. To reduce the computational time, Beesley et al. (Beesley et al., 2016) proposed approximations of the exact conditional distribution which are easier to handle in practice. However, assuming that $X$ and $Z$ are the same is quite restrictive, preventing us from fully exploiting the advantage of a mixture cure model in distinguishing between covariates with a curative or just a life-prolonging effect.
In this work, the ideas of Beesley et al. (Beesley et al., 2016) are extended to the more general model described in Section 2.1, considering three possible scenarios: the incomplete covariate belongs only to $X$, only to $Z$, or to both $X$ and $Z$. The scenario where the variable with missing values belongs to both $X$ and $Z$ is treated extensively in Section 3, while the cases where it belongs to only one of the two predictors (either the incidence or the latency component) are discussed in the appendices. Specifically, Appendix A deals with the case using the exact conditional distribution, while Appendix B addresses the case of the approximate conditional distribution. Further details about the method proposed by Beesley et al. (Beesley et al., 2016) and how it relates to this work are given in Section 3.
3 Methodology
In this section, we will describe our proposal for performing multiple imputation in cure models with incomplete covariates. As mentioned in Section 2.2, the main aspect of the multiple imputation procedure involves generating imputed values based on the conditional distribution of the incomplete variable given the other ones. Therefore, we will first derive the exact form of the conditional distribution and then outline the steps of the algorithm. Finally, we will introduce an approximation of the conditional distribution that enables us to generate imputed values at a reduced computational cost.
3.1 Using the exact conditional distribution
Let $W\in\mathbb{R}$ be the incomplete variable present in both the incidence and the latency model. The results for the cases where $W$ belongs either to the incidence or to the latency model are discussed in the Appendix. Since $W$ belongs to both $X$ and $Z$, there exist $1\leq j\leq p$ and $1\leq l\leq q$ such that $W=X^{(j)}=Z^{(l)}$. We denote the set of incidence and latency covariates excluding $W$ as $X^{(-j)}=X\setminus X^{(j)}$ and $Z^{(-l)}=Z\setminus Z^{(l)}$ respectively. To detail the proposed procedure in a more general context, we will consider two different choices for $W$, either a Gaussian or a Bernoulli random variable, conditional on the other covariates. However, results can be similarly generalized to other distributions.
Case 1. Let us first assume that $W|X^{(-j)},Z^{(-l)}$ follows a normal distribution, and that, at least approximately, the dependence of $W$ on $X^{(-j)}$ and $Z^{(-l)}$ is summarized by a change in the location (mean) parameter only. That is, we specify:
$W|X^{(-j)},Z^{(-l)}\sim\text{N}(\mu,\sigma),$
where
$\mu=\theta_{0}+\sum\limits_{b\neq j}^{p}\theta_{b}X_{b}+\sum\limits_{\begin{subarray}{c}s\neq l\\ Z_{s}\notin X\end{subarray}}^{d}{\theta_{p+s}}Z_{s}$ (6)
Note that we only consider variables in $Z$ that are not in $X$ ($Z_{s}\notin X$) to prevent any useless overlapping information. In this case, the complete data likelihood is given by
$\mathcal{L}(\alpha_{0},\alpha,\beta,H_{0},\mu,\sigma)$ $=\prod_{i=1}^{n}\left[\pi(X_{i})f_{u}(Y_{i}|Z_{i})\right]^{\Delta_{i}G_{i}}\left[\pi(X_{i})S_{u}(Y_{i}|Z_{i})\right]^{(1-\Delta_{i})G_{i}}$
×$\quad\times\left[1-\pi(X_{i})\right]^{(1-\Delta_{i})(1-G_{i})}f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)})$
$=\prod_{i=1}^{n}\left[\frac{e^{\alpha_{0}+\alpha^{T}X_{i}}}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}\left(h_{0}(Y_{i})e^{\beta^{T}Z_{i}}\right)^{\Delta_{i}}e^{-H_{0}(Y_{i})e^{\beta^{T}Z_{i}}}\right]^{G_{i}}$
×$\quad\times\left(\frac{1}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}\right)^{1-G_{i}}\left(e^{\frac{-(W_{i}-\mu_{i})^{2}}{2\sigma^{2}}}\right)$ (7)
Therefore, the exact distribution of $W$ given all the remaining covariates is as follows
$\begin{split}f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})&\propto\left(\frac{e^{\alpha_{0}+\alpha^{T}X_{i}}}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}e^{\beta^{T}Z_{i}\Delta_{i}}e^{-H_{0}(Y_{i})e^{\beta^{T}Z_{i}}}\right)^{G_{i}}\\ &\quad\times\left(\frac{1}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}\right)^{1-G_{i}}\left(e^{\frac{-(W_{i}-\mu_{i})^{2}}{2\sigma^{2}}}\right)\end{split}$ (8)
One can draw values to impute $W$ from this distribution using the Metropolis-Hasting algorithm (Metropolis et al., 1953; Hastings, 1970). This algorithm can draw values from a given generic probability distribution (which we will call the target distribution) by sampling from a Markov chain whose stationary distribution is the target distribution.
Case 2. Alternatively, $W|X^{(-j)},Z^{(-l)}$ could be a categorical random variable. For instance, if $W$ follows a Bernoulli distribution, we can specify:
$W|X^{(-j)},Z^{(-l)}\sim\text{Ber}(\operatorname{expit}(\mu)),$
where $\operatorname{expit}(x)=e^{x}/(1+e^{x})$ and $\mu$ is given by (6). The complete data likelihood is as in (3.1) where $f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)})$ is replaced by $P(W_{i}=w_{i}|X_{i}^{(-j)},Z_{i}^{(-l)})$ with $w_{i}\in\{0,1\}$. From it, and recalling that $\operatorname{logit}(x)=\log\bigl{(}\frac{x}{1-x}\bigr{)}$, the exact distribution of $W$ given all the remaining covariates takes the form
$\operatorname{logit}\left[P(W_{i}=1|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\right]=\log\left[\frac{P(W_{i}=1|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})}{P(W_{i}=0|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})}\right]$
$\qquad=G_{i}\Delta_{i}\beta_{l}-G_{i}H_{0}(Y_{i})e^{\sum_{s\neq l}^{d}{\beta_{s}Z_{i,s}}}(e^{\beta_{l}}-1)+G_{i}\alpha_{j}+\log\left(1+e^{\alpha_{0}+\sum_{b\neq j}^{p}{\alpha_{b}X_{i,b}}}\right)+$
$\qquad-\log\left(1+e^{\alpha_{0}+\alpha_{j}+\sum_{b\neq j}^{p}{\alpha_{b}X_{i,b}}}\right)+\mu_{i}$ (9)
Given that the conditional distributions depend on $\mu$ in both cases, one must draw values $\theta^{*}$ for the parameters $\theta$ before drawing values for the imputation of $W$. This is done by running a regression model (logistic regression when $W|X^{(-j)},Z^{(-l)}\sim$ Bernoulli and linear regression when $W|X^{(-j)},Z^{(-l)}\sim$ Normal), obtaining the point estimates $\hat{\theta}$ and the estimate for the covariance matrix $\hat{\Sigma}_{\theta}$ and drawing values $\theta^{*}$ from a Normal$\,(\hat{\theta},\hat{\Sigma}_{\theta})$ distribution.
3.2 Implementing multiple imputation in cure models
Let us consider the Cox PH cure model described in Section 2.1. To impute missing observations for any incomplete covariate by drawing values from its conditional distribution, we need to know $G$ and all the model parameters. Therefore, we use an iterative procedure where we first estimate the parameters, impute $G$, and then impute the missing covariate values. The pseudo-code of the corresponding algorithm is provided in Algorithm 1, and it can be briefly described as follows:
Algorithm 1 Imputation of missing covariates using exact conditional distribution
dataset of $n>0$ observations; number of iterations for chained equations $M>0$ ; number of imputed datasets $K>0$; $W|X^{(-j)},Z^{(-l)}\sim N(\mu,\sigma)$   or   $W|X^{(-j)},Z^{(-l)}\sim Ber(\operatorname{expit}(\mu))$
Imputed data for W
0: Initialization ($m=0$)
Assign values to $\alpha_{0}^{(0)},\alpha^{(0)},\beta^{(0)},H_{0}^{(0)}(t)$, and randomly fill the missing cases of $W$ producing $W^{(0)}$
for $m=1,\dots,M$ do
     1: Estimate $H_{0}(t)$
     for $i=1,\dots,n$ do
         a: Compute $\pi^{(m-1)}(X_{i})=\frac{\exp(\alpha_{0}^{(m-1)}+\alpha_{W}^{(m-1)}W_{i}^{(m-1)}+\alpha_{X}^{(m-1)T}X_{i}^{(-j)})}{1+\exp(\alpha_{0}^{(m-1)}+\alpha_{W}^{(m-1)}W_{i}^{(m-1)}+\alpha_{X}^{(m-1)T}X_{i}^{(-j)})}$
         b: Compute $S_{u}^{(m-1)}(t|Z_{i})=\exp\bigl{(}-H_{0i}^{(m-1)}(t)\exp(\beta_{W}^{(m-1)}W_{i}^{(m-1)}+\beta_{Z}^{(m-1)T}Z_{i}^{(-l)})\bigr{)}$
         c: Compute $q_{i}^{(m)}=\Delta_{i}+(1-\Delta_{i})\frac{\pi^{(m-1)}(X_{i})S_{u}^{(m-1)}(t|Z_{i})}{1-\pi^{(m-1)}(X_{i})+\pi^{(m-1)}(X_{i})S_{u}^{(m-1)}(t|Z_{i})}$
     end for
      Get $H_{0}^{(m)}(Y_{i})=\sum_{t_{j}<t_{i}}\frac{D(t_{j})}{\sum_{i\in R_{j}}q_{i}^{(m)}\exp(\beta_{W}^{(m-1)}W_{i}^{(m-1)}+\beta_{Z}^{(m-1)T}Z_{i}^{(-l)})}$
     2: Draw values for $\alpha_{0},\alpha,\beta$
      a: Compute $\hat{\alpha_{0}},\hat{\alpha}$ and $\hat{\Sigma}_{(\alpha_{0},\alpha)}$ from $\operatorname{logit}[P(G^{(m-1)}=1|W^{(m-1)},X^{(-j)})]$
      b: Draw $(\alpha_{0}^{(m)},\alpha^{(m)})$ from MVN$\bigl{(}(\hat{\alpha_{0}},\hat{\alpha}),\hat{\Sigma}_{(\alpha_{0},\alpha)}\bigr{)}$
      c: Compute $\hat{\beta}$ and $\hat{\Sigma}_{\beta}$ fitting $S_{u}(t|W^{(m-1)},Z^{(-l)})$ to the subjects such that $G_{i}=1$
      d: Draw $\beta^{(m)}$ from MVN$(\hat{\beta},\hat{\Sigma}_{\beta})$
     3: Impute $G$
     for all $G_{i}=$ NA  do draw from $\operatorname{logit}[P(G_{i}=1|W_{i}^{(m-1)},X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i}=0;\alpha_{0}^{(m)},\alpha^{(m)},\beta^{(m)})]$
     end for
     4: Impute W
     Specify $f(W_{i}|G_{i},\Delta_{i},Y_{i},X_{i}^{(-j)},Z_{i}^{(-l)};\Psi)$ and
     if $W\sim Ber(\operatorname{expit}(\mu))$ then
         a: Compute $\hat{\Psi}$ and $\hat{\Sigma}_{\Psi}$ from a logistic regression on $W_{i}^{(m-1)}$ using $X_{i}^{(-j)},Z_{i}^{(-l)}$
         b: Draw values $\Psi^{*}$ from MVN$(\hat{\Psi},\hat{\Sigma}_{\Psi})$
         c: for all $W_{i}^{(0)}=$ NA do draw imputed values from $\operatorname{logit}[P(W_{i}=1|G_{i},\Delta_{i},Y_{i},X_{i}^{(-j)},Z_{i}^{(-l)};\Psi^{*})]$
     else
         if $W\sim N(\mu,\sigma)$ then
              a: Compute $\hat{\Psi}$ and $\hat{\Sigma}_{\Psi}$ from a linear regression on $W_{i}^{(m-1)}$ using $X_{i}^{(-j)},Z_{i}^{(-l)}$
              b: Draw values $\Psi^{*}$ from MVN$(\hat{\Psi},\hat{\Sigma}_{\Psi})$
              c: for all $W_{i}^{(0)}=$ NA do Metropolis-Hastings draw using a random normal walk centered at $W_{i}^{(m-1)}$
         end if
     end if
     if $W\in\mathbb{R}^{a}$ with $a>1$ then
         repeat a, b and c for every $W$
     end if
end for
Repeat steps 1, 2, 3 and 4 to obtain $K$ complete datasets
  1. 0.
    The algorithm is initialized with values for $\alpha_{0}$, $\alpha$ and $\beta$ obtained by applying a Cox PH cure model to the complete-case dataset. By using a Breslow estimator (Breslow, 1974) on the same data, we obtain the initial values for $H_{0}(t)$. The missing cases of $W$ are randomly filled in.
  2. 1.
    The first step consists in estimating $H_{0}(Y)$, the baseline cumulative hazard rate for the individual times $(Y_{1},...,Y_{n})$ using the estimator in (5).
  3. 2.
    In the second step, we compute the estimates of the model parameters $\alpha_{0}$, $\alpha$ and $\beta$. For $\alpha_{0}$ and $\alpha$, this is done by fitting a logistic model with outcome $G$ and covariates $X$. Initially, the cure status $G$ has missing values for those censored units not in the right tail of the survival distribution (see step 3). In subsequent iterations, $G$ is filled in with the imputations obtained at the previous iteration. We then draw values $(\alpha_{0}^{*},\alpha^{*})$ from a multivariate Gaussian distribution, by using the asymptotic distribution of the ML estimates, e.g. MVN$\,((\hat{\alpha}_{0},\hat{\alpha}),\hat{\Sigma}_{(\alpha_{0},\alpha)})$. For $\beta$, this is done by fitting a Cox model to the observations where $G=1$, with follow-up times $Y$, event indicator $\Delta$, and covariates $Z$. We draw values $\beta^{*}$ from MVN$\,(\hat{\beta},\hat{\Sigma}_{\beta})$.
  4. 3.
    The following step consists of imputing the cure status $G$, which is known to be 1 for uncensored observations and assumed to be 0 for censored individuals with follow-up times after a certain cut-off point. As suggested in the literature (Sy and Taylor, 2000), we take the cut-off point equal to the last uncensored follow-up time. For censored observations until the cut-off point, $G$ is assumed to be missing at random conditional on $Y$ and $\Delta$ (Beesley et al., 2016), and it can be imputed using its exact conditional distribution
    $\operatorname{logit}\left[P(G_{i}=1|X_{i},Z_{i},Y_{i},\Delta_{i}=0)\right]=\log\left[\frac{S_{u}(Y_{i}|Z_{i})\pi(X_{i})}{1-\pi(X_{i})}\right]=-H_{0}(Y_{i})e^{\beta^{T}Z_{i}}+\left(\alpha_{0}+\alpha^{T}X_{i}\right)$
    which is derived from the likelihood introduced in (4). For the sake of clarity, let us recall that both $Z$ and $X$ include the variable with missing cases $W$.
  5. 4.
    Finally, we impute each missing covariate sequentially using the conditional distributions given in (8) or (3.1).
Each step makes use of the most recently imputed values of the incomplete variables and the most recently estimated model parameters.
3.3 Approximation using regression function
Drawing values from the exact conditional distributions in (8) and (3.1) can be very computationally intensive. Indeed, M-H algorithms are slowed down by the computation of complex target distributions and due to the slow convergence of such algorithms, one may need a very large number of iterations. Beesley et al. (Beesley et al., 2016) proposed an approximation and compared it to three alternatives from the literature. A simulation study indicated that the proposed approximation outperforms the others in terms of bias when the data has a cure fraction and $X=Z$. We extend this approach to cases where the sets of incidence and latency covariates differ. The following arguments will make use of the approximations
$\log(1+z)\approx\log(1+c)+\frac{z-c}{1+c}$ (10)
if $z$ is near $c$, and
$e^{aX+bY}\approx e^{a\bar{X}+b\bar{Y}}[1+a(X-\bar{X})+b(Y-\bar{Y})]$ (11)
if $Var(aX+bY)$ is small. The only constraint to consider for the approximation 10 to hold and be valid is that the argument of the logarithm function, $1+z$ and $1+c$, must be positive. Therefore, $z>-1$ and $c>-1$.
First, if we assume that $W|X^{(-j)},Z^{(-l)}\sim N(\mu,\sigma)$, then the $\log$ of the exact conditional distribution of $W$ is given by
$\log\left(f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\right)$ $=G_{i}\Delta_{i}\beta_{l}W_{i}-G_{i}H_{0}(Y_{i})e^{\beta^{T}Z_{i}}+G_{i}\alpha_{j}W_{i}-\log\left(1+e^{\alpha_{0}+\alpha^{T}X_{i}}\right)+$
$-\frac{1}{2\sigma^{2}}\left(W_{i}-\mu_{i}\right)^{2}+C$
where $C$ is a general additive term constant on $W_{i}$. Applying (10), (11), and a second-order Taylor approximation, one obtains
$\log\left(f(W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\right)$ $\approx-\frac{1}{2\sigma^{2}}W_{i}^{2}+\left[G_{i}\Delta_{i}\beta_{l}-G_{i}H_{0}(Y_{i})e^{\beta^{T}\bar{Z}}\beta_{l}+G_{i}\alpha_{j}+\right.$
$\quad\quad\left.-\frac{e^{\alpha_{0}+\alpha^{T}\bar{X}}}{1+e^{\alpha_{0}+\alpha^{T}\bar{X}}}\alpha_{j}+\frac{1}{\sigma^{2}}\mu_{i}\right]W_{i}+C$
Recalling that $\mu_{i}$ is a linear combination of $X_{i}^{(-j)}$ and $Z_{i}^{(-l)}$, one can see that the distribution of $W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i}$ is approximately Normal, where the mean is a linear combination of $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, $G_{i}$, $G_{i}\Delta_{i}$ and $G_{i}H_{0}(Y_{i})$. Therefore, a linear regression function with these covariates can be used to approximate the exact distribution of $W$.
Secondly, if we assume that $W|X^{(-j)},Z^{(-l)}\sim Ber(\operatorname{expit}(\mu))$, the conditional distribution of $W_{i}|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i}$ can be approximated using (10) and (11), obtaining
$\operatorname{logit}\left[P(W_{i}=1|X_{i}^{(-j)},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\right]$ $\approx G_{i}\Delta_{i}\beta_{l}-G_{i}H_{0}(Y_{i})(e^{\beta_{l}}-1)e^{\sum_{s\neq l}^{d}\beta_{s}\bar{Z}_{s}}\left[1+\sum_{s\neq l}^{d}\beta_{s}\left(Z_{i,s}-\bar{Z}_{s}\right)\right]$
$\quad+G_{i}\alpha_{j}+\frac{e^{\alpha_{0}+\sum_{b\neq j}^{p}\alpha_{b}\bar{X}_{b}}}{1+e^{\alpha_{0}+\sum_{b\neq j}^{p}\alpha_{b}\bar{X}_{b}}}\left[1+\sum_{b\neq j}^{p}\alpha_{b}\left(X_{i,b}-\bar{X}_{b}\right)\right]+$
$\quad-\frac{e^{\alpha_{0}+\alpha_{p}+\sum_{b\neq j}^{p}\alpha_{b}\bar{X}_{b}}}{1+e^{\alpha_{0}+\alpha_{p}+\sum_{b\neq j}^{p}\alpha_{b}\bar{X}_{b}}}\left[1+\sum_{b\neq j}^{p}\alpha_{b}\left(X_{i,b}-\bar{X}_{b}\right)\right]+\mu_{i}+C$
This is a linear combination of $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, $G_{i}$, $G_{i}\Delta_{i}$, $G_{i}H_{0}(Y_{i})$ and $G_{i}H_{0}(Y_{i})Z_{i}^{(-l)}$. Therefore, the exact distribution of $W$ can be approximated using a logistic regression function with these covariates as the last step of the imputation procedure presented in Section 3.2.
Details about the scenarios where the variable with missing values $W$ belongs only to the incidence or latency model are provided in Appendix B. If $W$ belongs only to the incidence, the conditional distribution of $W$ can be approximated by linear or logistic regression with covariates $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, and $G_{i}$. When $W$ belongs only to the latency, the conditional distribution of W can be approximated by linear or logistic regression with covariates $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, $G_{i}\Delta_{i}$, $G_{i}H_{0}(Y_{i})$, and (for Case 2) $G_{i}H_{0}(Y_{i})Z_{i}^{(-l)}$.
4 Simulation
4.1 Simulation design
We conducted a simulation study to compare the performance of the imputation methods for cure models presented in Section 3. The comparison was carried out considering several scenarios, defined as a function of the missing data percentages, model parameters, and variables’ distributions.
In this simulation study, we generated 1000 datasets, each containing 500 observations. The variables used in the study are $W$, $X$ and $Z$. Both $X$ and $Z$ follow a Bernoulli distribution with a probability of 0.5, while $W$ is either $Ber(0.5)$ or $N(0.5,1)$. The choice of a Normal variable with a mean of 0.5 rather than a standard Normal was made to ensure the stability of the cure rate when varying W’s distribution.
Variable $X$ is a covariate in the incidence model, $Z$ is a covariate in the latency model, and both of them have no missing values. Variable $W$ is a covariate in the models for both the incidence and the latency, and it is either missing completely at random (MCAR) with 15% or 30% missing values or missing at random (MAR) with 30% missing observations. To produce missing cases following a MAR mechanism, we used a multivariate amputation procedure (Schouten et al., 2018) in which the probability of $W$ to be missing depends on both $X$ and $Z$.
The time to event $T$ is generated from a Weibull PH model of the form
$S(t)=\exp{(-\lambda t^{\rho}\exp{(\beta_{1}W+\beta_{2}Z)})}$
with $\lambda=0.25$, $\rho=1.45$, and truncated at 8, to satisfy the sufficient follow-up assumption. For cured individuals (having $G=0$), $T$ is set to an extremely large value, as they never experience the event of interest.
The time to censoring $C$ is drawn from an exponential distribution with a parameter equal to 0.08 when $W$ is Bernoulli distributed, and to 0.1 when $W$ is Normally distributed. Both censoring distributions are truncated at 10.
The complete uncured indicator $G$ is generated as a Bernoulli random variable with probability $\pi$ defined by a logistic regression of the form
$P(G=1)=\frac{\exp{(\alpha_{0}+\alpha_{1}W+\alpha_{2}X)}}{1+\exp{(\alpha_{0}+\alpha_{1}W+\alpha_{2}X)}}$ (12)
In practice, we only use the complete indicator drawn from (12) for generating the follow-up times $Y$, namely to assign large values of $Y$ to cured individuals. Instead, for multiple imputation in the cure model, we use the observed uncured indicator, which is 1 for uncensored observations, 0 for observations in the plateau, and unknown otherwise.
We consider two sets of parameter values for the incidence and latency models. The first one is $\alpha_{0}=1$, $\alpha_{1}=-1$, $\alpha_{2}=0.5$, $\beta_{1}=-0.2$ and $\beta_{2}=0$. This set of parameters is based on the motivating application and therefore is only used when $W$ is Bernoulli distributed with around 15% of observations missing completely at random. This scenario leads to an average cure rate of $\sim 33$%, an average censoring rate of $\sim 45$%, and an average $\sim 18$% observations in the plateau.
The second set of parameters is $\alpha_{0}=0.1$ and $\alpha_{1}=\alpha_{2}=\beta_{1}=\beta_{2}=0.5$. These values are chosen to illustrate a more general setting. For this choice of parameters, we consider $W$ as Normally distributed with around 30% of observations missing at random or missing completely at random. Both the scenarios in which the missing mechanism is MCAR and MAR lead to an average cure rate of $\sim 36$%, an average censoring rate of $\sim 46$%, and an average $\sim 17$% observations in the plateau.
To summarize this simulation design, the three considered scenarios are the following:
  • Scenario A: MCAR data with 15% missingness; $W\sim$ Bernoulli; $\alpha_{0}=1$, $\alpha_{1}=-1$, $\alpha_{2}=0.5$, $\beta_{1}=-0.2$ and $\beta_{2}=0$.
  • Scenario B: MCAR data with 30% missingness; $W\sim$ Normal; $\alpha_{0}=0.1$ and $\alpha_{1}=\alpha_{2}=\beta_{1}=\beta_{2}=0.5$.
  • Scenario C: MAR data with 30% missingness; $W\sim$ Normal; $\alpha_{0}=0.1$ and $\alpha_{1}=\alpha_{2}=\beta_{1}=\beta_{2}=0.5$.
For the multiple imputation procedure using the exact conditional distribution in the case of Gaussian $W$, we use a Normal distribution with unit variance in the Metropolis-Hastings sampling algorithm. The sampler runs for 500 iterations before recording results with 100 iterations between successive samples.
For the multiple imputation procedure using the approximate conditional distributions, we use the mice (van Buuren and Groothuis-Oudshoorn, 2011) R package. For both the exact and the approximated method, we set the number of iterations of the chained-equations to 10 (Pan and Wei, 2016) and the number of imputed datasets to 10.
There are two packages to estimate cure models in R, smcure (Cai et al., 2012) and curephEM (Hou et al., 2018). We used the latter with the default values for the convergence criteria and bootstrap sampling for standard error estimation.
4.2 Simulation results
In Table 1, we present the results of simulations for the three different scenarios.
Table 1: Simulation study results. Evaluation metrics, $B=1000$ samples
      MSE1 (CI width2) Coverage3
Method $\alpha_{0}$ $\alpha_{1}$ $\alpha_{2}$ $\beta_{1}$ $\beta_{2}$
Scenario A
Full data 0.04 (0.80) 0.96 0.05 (0.92) 0.95 0.05 (0.90) 0.96 0.02 (0.53) 0.95 0.02 (0.51) 0.96
Complete-case 0.05 (0.87) 0.95 0.06 (1.00) 0.96 0.06 (0.98) 0.96 0.02 (0.58) 0.95 0.02 (0.56) 0.95
Exact 0.04 (0.83) 0.96 0.06 (1.00) 0.97 0.05 (0.91) 0.96 0.02 (0.57) 0.96 0.02 (0.51) 0.96
Approximate 0.04 (0.84) 0.95 0.06 (1.00) 0.96 0.05 (0.91) 0.96 0.02 (0.57) 0.95 0.02 (0.51) 0.96
Scenario B
Full data 0.03 (0.63) 0.94 0.01 (0.46) 0.94 0.05 (0.86) 0.96 0.01 (0.29) 0.94 0.02 (0.53) 0.95
Complete-case 0.04 (0.75) 0.95 0.02 (0.55) 0.96 0.07 (1.04) 0.96 0.01 (0.35) 0.94 0.03 (0.63) 0.94
Exact 0.03 (0.65) 0.94 0.01 (0.60) 0.99 0.04 (0.86) 0.97 0.01 (0.35) 0.94 0.02 (0.54) 0.94
Approximate 0.03 (0.66) 0.95 0.02 (0.60) 0.96 0.06 (0.89) 0.94 0.01 (0.35) 0.94 0.02 (0.56) 0.95
Scenario C
Full data 0.03 (0.63) 0.95 0.01 (0.46) 0.94 0.05 (0.86) 0.97 0.01 (0.29) 0.96 0.02 (0.53) 0.94
Complete-case 0.35 (0.87) 0.32 0.03 (0.68) 0.93 0.12 (1.28) 0.94 0.01 (0.33) 0.96 0.03 (0.60) 0.93
Exact 0.03 (0.65) 0.94 0.01 (0.60) 0.99 0.04 (0.86) 0.97 0.01 (0.34) 0.93 0.02 (0.54) 0.94
Approximate 0.03 (0.68) 0.95 0.03 (0.79) 0.97 0.05 (0.91) 0.96 0.01 (0.35) 0.93 0.02 (0.55) 0.95
1 MSE: mean squared error
2 Width of 95% confidence intervals
3 Empirical coverage of 95% confidence intervals
The performance of the imputation methods was assessed using different metrics, specifically the mean squared error (MSE), the width of 95% confidence intervals (CI width), and the empirical coverage of 95% confidence intervals (Coverage). The full data scenario provides a baseline for comparison, as it describes these metrics for models fitted to the original data, without missing values.
In the scenario with 15% MCAR (Scenario A), the complete-case analysis, obtained by discarding units with missing values for $W$ and then fitting a cure model on the complete data only, exhibits slightly higher MSE and wider confidence intervals compared to the full data scenario. As expected, estimates from the complete case analysis are not biased but have a larger variance since fewer observations are being used. Both the exact and the approximate imputation methods performed similarly to the full data scenario, with only minor differences in CI widths, only for the estimates of the effects associated with $W$. In particular, we observe that multiple imputation lead to decreased variance for the estimates of the parameters corresponding to fully observed variables.
When the missingness is increased to 30%, still under an MCAR mechanism (Scenario B), this trend is confirmed: the complete-case analysis shows higher MSE and wider confidence intervals, while the imputation methods (via either the exact or approximate procedure) yield results close to the full data scenario. Although there are slight variations, these methods effectively preserve empirical coverage, with a slightly greater CI length when the effects of $W$ are considered.
These results show that the proposed approaches provide reliable parameter estimates even in the presence of a moderate to high proportion of missing data. However, the true strength of our methods becomes evident in the MAR scenario (Scenario C). Here, the proposed imputation methods significantly outperform the complete-case analysis according to all evaluation metrics. This demonstrates that when data are missing at random, both exact and approximate imputation methods can effectively mitigate the impact of missing data, producing estimates close to those obtained with full data. In contrast, the complete-case analysis performs poorly, exhibiting a marked increase in MSE and a decrease in coverage as a result of the bias. Specifically, for $\alpha_{0}$, the MSE of the complete-case approach increased to 0.35 and the coverage decreased to 0.32. Additionally, the CI width for $\alpha_{2}$ was considerably wider than that obtained with the imputation methods. Moreover, the estimate $\beta_{2}$ shows a slight deterioration compared to scenarios A and B. Indeed, when using complete-case analysis, the missing data of $W$ leads to the removal of observations that might be critical for accurately estimating $\alpha_{2}$ and $\beta_{2}$, especially if $W$ is correlated with $X$ and $Z$. This results in biased estimates and an increase in the MSE. In other words, by eliminating observations with missing $W$ data, the model’s ability to capture the true relationships involving $X$ and $Z$ may be reduced.
The exact conditional distribution method outperforms the approximate method in terms of CI width, indicating more precise estimates.
Across all scenarios, both the exact and approximate methods consistently provide reliable estimates, maintaining the empirical coverage of the 95% confidence intervals. Particularly in cases where covariate $W$ is missing at random, the proposed methods significantly outperform the complete-case approach, demonstrating their ability to effectively handle complex missing data patterns.
4.3 Algorithm performance under model misspecification
Appendix A and B provide the exact and approximate distributions for cases where the variable $W$ is included in only the incidence or the latency of the Cox PH cure model. However, we may lack information about the impact of $W$ on the two components. Therefore, we aim to study the performance of the proposed imputation algorithm when the model is misspecified, e.g. when $W$ is present in only one of the two components and we assume it is in both. To do this, we conduct a simulation study, examining two scenarios:
  • Scenario D: MAR data with 30% missingness; $W\sim$ Normal; $\alpha_{0}=0.1$, $\alpha_{1}=\alpha_{2}=\beta_{2}=0.5$, and $\beta_{1}=0$
  • Scenario E: MAR data with 30% missingness; $W\sim$ Normal; $\alpha_{0}=0.1$, $\alpha_{2}=\beta_{1}=\beta_{2}=0.5$, and $\alpha_{1}=0$
These scenarios are similar to Scenario C introduced in section 4.1, but they differ in that the coefficient associated with the missing covariate $W$ is set to zero in either the latency component (Scenario D, $\beta_{1}=0$) or the incidence component (Scenario E, $\alpha_{1}=0$). This allows us to test the algorithm’s ability to handle situations where the missing variable affects only one component of the Cox PH cure model.
Importantly, our focus is testing the algorithm’s performance under the MAR scenario. We expect that if the algorithm performs well in these more complex conditions, it will likely do so in scenarios where the data is MCAR or the percentage of missing data is lower.
This simulation study provides insights into the robustness of the mixture cure model across the two distinct scenarios. If the imputation algorithm performs well under these misspecifications, we can derive that our model is robust and does not require prior knowledge of the covariate’s precise impact on incidence or latency. Instead, the model, through its fit to the available data, can suggest the appropriate placement of different covariates based on the significance of the coefficient estimates.
In Table 2, simulation results for scenarios E and D are shown. Analogously to what is seen in Table 1, the performance of the imputation methods was assessed using the MSE, the CI width, and the empirical coverage of 95% confidence intervals.
Table 2: Simulation results under model misspecification. Evaluation metrics, $B=1000$ samples
      MSE1 (CI width2) Coverage3
Method $\alpha_{0}$ $\alpha_{1}$ $\alpha_{2}$ $\beta_{1}$ $\beta_{2}$
Scenario D
Full data 0.03 (0.63) 0.95 0.01 (0.47) 0.96 0.05 (0.88) 0.94 0.00 (0.28) 0.97 0.02 (0.54) 0.94
Complete-case 0.33 (0.86) 0.31 0.03 (0.69) 0.96 0.15 (1.31) 0.93 0.01 (0.32) 0.95 0.02 (0.62) 0.95
Exact 0.03 (0.65) 0.94 0.03 (0.58) 0.91 0.05 (0.86) 0.94 0.01 (0.33) 0.98 0.02 (0.54) 0.94
Approximate 0.03 (0.69) 0.96 0.03 (0.87) 0.98 0.05 (0.92) 0.93 0.01 (0.34) 0.97 0.02 (0.54) 0.94
Scenario E
Full data 0.02 (0.61) 0.96 0.01 (0.41) 0.94 0.04 (0.81) 0.95 0.01 (0.30) 0.96 0.02 (0.55) 0.96
Complete-case 0.31 (0.83) 0.33 0.02 (0.58) 0.95 0.10 (1.14) 0.91 0.01 (0.34) 0.95 0.03 (0.63) 0.94
Exact 0.03 (0.62) 0.95 0.03 (0.53) 0.92 0.04 (0.82) 0.95 0.01 (0.35) 0.90 0.02 (0.56) 0.95
Approximate 0.03 (0.70) 0.96 0.03 (0.72) 0.95 0.04 (0.83) 0.95 0.01 (0.36) 0.94 0.02 (0.58) 0.96
1 MSE: mean squared error
2 Width of 95% confidence intervals
3 Empirical coverage of 95% confidence intervals
As before, the data without missing observations serves as a benchmark.
Quite similar to results in Table 1, the complete-case analysis shows a significant increase in MSE for $\alpha_{0}$, reaching 0.33 in Scenario D, with a corresponding coverage of only 0.31. This result reflects the method’s vulnerability to bias when missing data is not appropriately addressed. The complete-case method also displays increased CI widths, especially for $\alpha_{2}$, highlighting the loss of information and precision due to missing data.
The exact method generally maintains low MSE values in both scenarios and good coverage for most parameters. However, its performance sometimes falls short of the approximate method, as seen with $\alpha_{1}$ in Scenario D, where the exact method yields a coverage of 0.91 compared to 0.98 of the approximate method. The improved coverage of the approximate method can be attributed to its increased CI width, which provides a more conservative estimate, thereby encompassing more potential values of the parameter. The exact method’s performance suggests it may be more sensitive to model assumptions, leading to increased bias when these assumptions are not met. Despite this, the results support conclusions about the general robustness of the model. The slightly poorer performance compared to the approximate method could be due to the tighter linkage between the missing data variable $W$ and the linear predictors, resulting in higher sensitivity to model assumptions and less flexibility in handling deviations.
The approximate method shows consistent performance with low MSEs in both scenarios and high coverage, highlighting its adaptability in dealing with model deviations. It achieves slightly better coverage for parameters like $\beta_{1}$ in Scenario E, reaching 0.94 compared to the exact method’s 0.90.
In Scenario E, the complete-case analysis exhibits substantial MSE values for $\alpha_{0}$ and poor coverage, confirming its inadequacy in scenarios with high levels of missingness. Conversely, the approximate method maintains reliable coverage and CI widths across parameters due to its adaptability and reduced reliance on strict assumptions.
Overall, these results suggest that the approximate method is particularly well-suited for real-world applications where knowledge of model structure is limited and data are affected by several sources of error. Its robust performance across scenarios D and E reinforces the idea that it is a viable alternative to the exact method.
5 Case study
Our method is applied to data from the MRC BO06/EORTC 80931 randomised controlled trial (RCT) for patients with localised resectable high-grade osteosarcoma. Osteosarcoma is a malignant bone tumour which mostly affects children, adolescents and young adults. Patients with osteosarcoma cannot be successfully treated with surgery alone, but adjuvant chemotherapy has been proven to significantly increase survival (Link et al., 1986; Eilber et al., 1987). The impact of doses and dose intensities is a topic of continuous discussion in this context (Lancia et al., 2019) and it is also relevant to other cancer types. Several previous studies have identified the presence of long-term disease-free survivors among osteosarcoma patients, who are practically considered as ‘cured’. Indeed, less than five per cent of patients experience a relapse after five or more years of follow-up. This implies that a patient falls into one of two groups once the main treatment has been administered: those who have been healed and those who have not, meaning they will eventually suffer from the progression of their disease.
The BO06 clinical trial includes 497 patients diagnosed between May 1993 and September 2002 (Lewis et al., 2007), and randomly assigned to receive either a conventional two-drug regimen (Reg-C), consisting of six 3-week cycles of doxorubicin (DOX, 75 mg/m2) and cisplatin (CDDP, 100 mg/m2) or a dose-intensified (DI) regimen (Reg-DI), consisting of the same doses administered twice a week and supported by G-CSF (5µg/kg daily).
The aim of the MRC BO06/EORTC 80931 clinical trial was to investigate whether increasing the DI would improve the survival of patients with nonmetastatic limb osteosarcoma. The primary outcome measure is progression-free survival (PFS) since the end of treatment, as a cure could occur at any time during the treatment.
We excluded from the study the patients who did not receive chemotherapy, had abnormal dosages of one or both agents (more than 1.25×prescribed dose), did not have surgery, died or experienced disease progression during treatment (these last two groups were excluded because the primary outcome is progression-free survival calculated from the end of treatment), resulting in a sample of 429 individuals.
Some of the patients’ baseline characteristics are shown in Table 3, more details about the patients and chemotherapy can be found in the primary publication of the trial (Lewis et al., 2007).
Table 3: Baseline characteristics of patients in the BO06 trial.
Characteristic $N=429$
Allocated treatment, n (%)
   Reg-C 207 (48%)
   Reg-DI 222 (52%)
Sex, n (%)
   Female 173 (40 %)
   Male 256 60%)
Age, median (range) 15.2 (3.5, 40.7)
Histological response, n (%)
   Poor 214 (56%)
   Good 165 (44%)
   Unknown 50
Previous study (Musta et al., 2022) showed that the histologic response is strongly prognostic for the cure fraction, while for the uncured patients, there is no clear indication that good histologic response is associated with longer PFS. On the other hand, there is no evidence of a statistically significant effect of the allocated regimen on the cure fraction or the survival rates, suggesting that the intensified treatment might not be associated with better cure chances and longer PFS.
We fit a Cox proportional hazards cure model to the complete data using the variables histological response and sex as predictors in the incidence part and histological response and assigned treatment as predictors in the latency part of the mixture cure model. Details regarding the test for the sufficient follow-up assumption are provided in Appendix C. We performed multiple imputation of the histological response using both the exact and the approximated approaches, as this variable suffers from $\sim 12$% missing values. Figure 2 illustrates the results of the Cox proportional hazards cure model for time-to-osteosarcoma recurrence.
(a) Complete-case analysis
(b) Multiple imputation using the exact conditional distribution
(c) Multiple imputation using the approximate conditional distribution
Figure 2: Cox proportional hazards cure model for time-to osteosarcoma recurrence. Odds Ratios for the incidence component and Hazard Ratios for the latency component of the mixture cure models, together with 95% confidence intervals.
This figure displays the odds ratios for the incidence component and the hazard ratios for the latency component of the mixture cure models, along with 95% confidence intervals. Panel (a) presents the results of the complete-case analysis, while panels (b) and (c) show the outcomes of multiple imputation using the exact and approximate conditional distributions, respectively.
Point estimates and confidence intervals are very similar between the two imputation approaches. Based on the simulation results, we do not expect big differences between the analyses based on imputed data and the complete-case one, as the percentage of missingness in the histological response variable is relatively small.
The only minor difference is that confidence intervals tend to be narrower for the imputation approaches than for complete-case analysis. The most notable difference between the imputation and complete-case fits is in the estimates of the model intercept, mainly in the width of his confidence interval, which is much wider in the complete-case setting.
6 Discussion
Our work on imputation methods for mixture cure models in the presence of missing covariate values emphasizes the importance of flexible imputation techniques. In this paper, we focus on improving the methodology for imputing partially observed covariates by accommodating potentially distinct sets of covariates for the cure probability and the survival of uncured patients.
Results show that both exact and approximate imputation approaches produce similar results. Notably, the imputation methods yield narrower confidence intervals compared to the complete-case analysis, enhancing the precision of our estimates.
The simulation results suggest that the proposed methods are robust to the presence of missing values, particularly under the Missing at Random (MAR) scenario, where they outperform the complete-case analysis. This highlights the effectiveness of multiple imputation in handling complex missing data patterns, preserving empirical coverage, and reducing Mean Squared Error (MSE). The exact imputation method slightly outperforms the approximate method in terms of confidence interval width, yielding more precise estimates. However, when the model is misspecified, i.e. the covariate with missing values is wrongly included in both latency and incidence components, the approximate method has better coverage probabilities. These findings also point to the advantage of differentiating covariates between incidence and latency, providing a more nuanced and accurate analysis.
For the BO06 clinical trial data on osteosarcoma, the application of the Cox proportional hazards cure model reveals several key findings. First, histological response emerges as a strong prognostic factor for the cure fraction, indicating that patients who respond well to initial treatment are more likely to be cured. However, again there is no clear indication that a good histological response is associated with longer progression-free survival (PFS) for uncured patients. This underscores the importance of having the possibility of differentiating between factors that influence cure probability from those that affect survival among uncured patients.
Overall, the enhanced methodology for multiple imputation, accommodating different sets of covariates for incidence and latency, proves to be a valuable tool in the analysis of survival data with missing covariates. This approach not only improves the accuracy and reliability of the estimates but also provides a more nuanced understanding of the factors influencing both cure probability and survival of uncured patients in osteosarcoma. The flexibility in covariate selection allows for more tailored and precise modelling, ultimately enhancing the interpretability and applicability of the results in clinical practice.
As a further development of this work, one might consider an alternative way of performing multiple imputation based on the observed likelihood (instead of the complete one), which avoids considering the latent cure status as an additional variable with missing values. When there is only one missing variable, this would avoid the chained equations since there would be no need to impute G. However, such an approach would require the estimation of the parameters $\alpha$ and $\beta$ in Step 2 of Algorithm 1 by fitting a mixture cure model and the variance needs to be estimated via computationally intensive bootstrap. Hence, we did not explore this method further in this work.
Supplementary material
The source codes for implementing the imputation algorithm and for reproducing the simulations are available at https://github.com/martacip/mi_curemodels.
Acknowledgments
MRC Clinical Trials Unit at UCL (London, UK) and the European Organisation for Research and Treatment of Cancer (EORTC) are gratefully acknowledged for making data from the MRC BO06/EORTC 80931 trial available for this analysis. The trial was supported and coordinated by the U.K. MRC and the EORTC.
Appendix A - Imputation using the exact conditional distribution
We derive an exact imputation model for imputing both normal and binary covariates in the case in which the partially observed covariate belongs either to the incidence component or to the latency component of the Cox PH mixture cure model.
A.1 Incomplete covariate only in the incidence
Case 1. Suppose $W|X^{(-j)},Z\sim Normal(\mu,\sigma)$ where
$\mu=\theta_{0}+\sum\limits_{b\neq j}^{p}\theta_{b}X_{b}+\sum\limits_{\begin{subarray}{c}s\\ Z_{s}\notin X\end{subarray}}^{d}{\theta_{p+s}}Z_{s}$ (13)
Updating the complete data likelihood for the Cox PH cure model given in (3.1), we obtain the following exact conditional distribution for $W$
$\begin{split}f(W_{i}|X_{i}^{(-j)},Z_{i},Y_{i},\Delta_{i},G_{i})\propto\left(\frac{e^{\alpha_{0}+\alpha^{T}X_{i}}}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}\right)^{G_{i}}\left(\frac{1}{1+e^{\alpha_{0}+\alpha^{T}X_{i}}}\right)^{1-G_{i}}\left(e^{\frac{-(W_{i}-\mu_{i})^{2}}{2\sigma^{2}}}\right)\end{split}$ (14)
Case 2. Let assume that $W|X^{(-j)},Z\sim Bernoulli(\operatorname{expit}(\mu))$, where $\mu$ is given by (13), then the exact distribution of $W$ takes the form
$\operatorname{logit}\bigl{[}P(W_{i}=1|X_{i}^{(-j)},Z_{i},Y_{i},\Delta_{i},G_{i})\bigr{]}=G_{i}\alpha_{j}+\log\bigl{(}1+e^{\alpha_{0}+\sum_{b\neq j}^{p}{\alpha_{b}X_{i,b}}}\bigr{)}-\log\bigl{(}1+e^{\alpha_{0}+\alpha_{j}+\sum_{b\neq j}^{p}{\alpha_{b}X_{i,b}}}\bigr{)}+\mu_{i}$ (15)
A.2 Incomplete covariate only in the latency
Case 1. If we assume that $W|X,Z^{(-l)}\sim\text{Normal}(\mu,\sigma)$ and we define $\mu$ as
$\mu=\theta_{0}+\sum\limits_{b}^{p}\theta_{b}X_{b}+\sum\limits_{\begin{subarray}{c}s\neq l\\ Z_{s}\notin X\end{subarray}}^{d}{\theta_{p+s}}Z_{s}$ (16)
we can derive the following exact conditional distribution for $W$
$\begin{split}f(W_{i}|X_{i},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\propto\left(e^{\beta^{T}Z_{i}\Delta_{i}}e^{-H_{0}(Y_{i})e^{\beta^{T}Z_{i}}}\right)^{G_{i}}\left(e^{\frac{-(W_{i}-\mu_{i})^{2}}{2\sigma^{2}}}\right)\end{split}$ (17)
Case 2. Suppose that $W|X,Z^{(-l)}\sim\text{Bernoulli}(\operatorname{expit}(\mu))$, where $\mu$ is as in (16), then $W$ follows the distribution
$\operatorname{logit}\bigl{[}P(W_{i}=1|X_{i},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\bigr{]}=G_{i}\Delta_{i}\beta_{l}-G_{i}H_{0}(Y_{i})e^{\sum_{s\neq l}^{d}{\beta_{s}Z_{i,s}}}(e^{\beta_{l}}-1)+\mu_{i}$ (18)
Appendix B - Imputation using the approximate conditional distribution
Similarly to what was discussed in Section 3.3 and by utilizing approximations (10) and (11), we can derive an approximate version of the exact distributions shown in Appendix A, which allows us to impute the missing values of W at a reduced computational cost.
B.1 Incomplete covariate only in the incidence
Case 1. We assume that $W|X^{(-j)},Z\sim Normal(\mu,\sigma)$ and that $\mu$ is equal to 13. An approximation of the exact conditional distribution in 14 is given by
$\log\left(f(W_{i}|X_{i}^{(-j)},Z_{i},Y_{i},\Delta_{i},G_{i})\right)\approx-\frac{1}{2\sigma^{2}}W_{i}^{2}+\left[G_{i}\alpha_{j}-\frac{e^{\alpha_{0}+\alpha^{T}\bar{X}}}{1+e^{\alpha_{0}+\alpha^{T}\bar{X}}}\alpha_{j}+\frac{1}{\sigma^{2}}\mu_{i}\right]W_{i}+C$ (19)
where $C$ is a constant term. A linear regression function with the covariates $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, and $G_{i}$ can be used to approximate the exact distribution of $W$.
Case 2. If $W|X^{(-j)},Z$ is Bernoulli distributed with parameter $\operatorname{expit}(\mu)$, with $\mu$ equal to 13, an approximate imputation model is as follows
$\operatorname{logit}\left[P(W_{i}=1|X_{i}^{(-j)},Z_{i},Y_{i},\Delta_{i},G_{i})\right]$ $\approx G_{i}\alpha_{j}+\frac{e^{\alpha_{0}+\sum_{b\neq j}^{p}{\alpha_{b}\bar{X}_{b}}}}{1+e^{\alpha_{0}+\sum_{b\neq j}^{p}{\alpha_{b}\bar{X}_{b}}}}\left[1+\sum_{b\neq j}^{p}{\alpha_{b}\left(X_{i,b}-\bar{X}_{b}\right)}\right]+$
$-\frac{e^{\alpha_{0}+\alpha_{j}+\sum_{b\neq j}^{p}{\alpha_{b}\bar{X}_{b}}}}{1+e^{\alpha_{0}+\alpha_{j}+\sum_{b\neq j}^{p}{\alpha_{b}\bar{X}_{b}}}}\left[1+\sum_{b\neq j}^{p}{\alpha_{b}\left(X_{i,b}-\bar{X}_{b}\right)}\right]+\mu_{i}+C$ (20)
A logistic regression function with the covariates $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, and $G_{i}$ can be used to approximate the exact distribution of $W$.
B.2 Incomplete covariate only in the latency
Case 1. We can approximate the above equation (17) by
$\log\left(f(W_{i}|X_{i},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\right)\approx-\frac{1}{2\sigma^{2}}W_{i}^{2}+\left[G_{i}\Delta_{i}\beta_{l}-G_{i}H_{0}(Y_{i})e^{\beta^{T}\bar{Z}}\beta_{l}+\frac{1}{\sigma^{2}}\mu_{i}\right]W_{i}+C$ (21)
A linear regression function with the covariates $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, $G_{i}\Delta_{i}$, and $G_{i}H_{0}(Y_{i})$ can be used to approximate the exact distribution of $W$.
Case 2. A simpler approximated model for (18) is given by
$\operatorname{logit}\left[P(W_{i}=1|X_{i},Z_{i}^{(-l)},Y_{i},\Delta_{i},G_{i})\right]$ $\approx G_{i}\Delta_{i}\beta_{l}-G_{i}H_{0}(Y_{i})(e^{\beta_{l}}-1)e^{\sum_{s\neq l}^{d}{\beta_{s}\bar{Z}_{s}}}\left[1+\sum_{s\neq l}^{d}{\beta_{s}\left(Z_{i,s}-\bar{Z}_{s}\right)}\right]$
$\qquad+\mu_{i}+C$ (22)
A logistic regression function with the covariates $X_{i}^{(-j)}$, $Z_{i}^{(-l)}$, $G_{i}\Delta_{i}$, $G_{i}H_{0}(Y_{i})$, and $G_{i}H_{0}(Y_{i})Z_{i}^{(-l)}$ can be used to approximate the exact distribution of $W$.
Appendix C - Testing the sufficient follow-up assumption for the osteosarcoma data
We test the sufficient follow-up assumption for the osteosarcoma data analyzed in Section 5. Currently there exist no test uniformly over all covariate values so we apply the tests proposed by Xie et al. [Xie et al., 2023] and Yuen and Musta [Yuen and Musta, 2024] separately for patients with good and poor histologic response, since that is the main prognostic factor for the cure probability.
The test proposed by Xie et al. [Xie et al., 2023] considers the null hypothesis of sufficient follow-up $H_{0}:\tau_{F_{u}}\leq\tau_{C}$, where $\tau_{F_{u}}$ and $\tau_{C}$ denote, respectively, the right end points of the distribution of the event times for the uncured patients and of the censoring times. For patients with good histologic response, such test does not reject the null hypothesis of sufficient follow-up since $2(Y_{(n)}-\tilde{Y}_{(m)})<Y_{(n)}$, where $Y_{(n)}$ denotes the largest observed time and $\tilde{Y}_{(m)}$ denotes the largest observed event time. For patients with poor histologic response, the p-value of the test is 0.038 meaning that the null hypothesis is rejected at level 0.05.
However, it is argued by Yuen and Musta [Yuen and Musta, 2024] that one could instead test a more relaxed notion of practically sufficient follow-up. The null hypothesis is that the follow-up is insufficient $\tilde{H}_{0}:\tau_{C}\leq q_{1-\epsilon}$, where $q_{1-\epsilon}$ denotes the $(1-\epsilon)$-quantile of $F_{u}$ and $\epsilon$ is a small number chosen by the user. It essentially tests whether there is less that $\epsilon$ probability for the event to happen after the end of the study. In practice this would be considered as a sufficiently long follow-up if $\epsilon$ is around $1$-$2\%$. The test requires also the choice of a parameter $\tau$ such that the chances of the event happening after $\tau$ are negligible, i.e. a study of length $\tau$ would have been certainly sufficient. We apply the test for two different choices of $\epsilon$ and $\tau$ and p-values are given in Table 4. The results indicate that the follow-up is clearly sufficient for patients with good histologic response and borderline sufficient for those with poor histologic response.
good hist. poor hist.
$\tau$ $\tau$
$\epsilon$ 15 years 20 years 15 years 20 years
0.01 0.000 0.000 0.095 0.131
0.02 0.000 0.000 0.046 0.093
Table 4: P-values for the testing the null hypothesis $\tilde{H}_{0}$ of practically sufficient follow-up for patients with good and poor histologic response.